A lossless transmission line is used to deliver an AC signal from Liverpool to Wirral. The length of the transmission line is 6 km. The phase velocity is the same as in free space. The transmission line is terminated by a load at the Wirral side with an impedance of 100 Ω. j) If the characteristic impedance of the transmission line is 100 Ω and the AC b) signal has a frequency of 50 Hz, what is the input impedance seen at the Liverpool side? ii) If the characteristic impedance of the transmission line is 20 Ω and the AC signal has a frequency of 50 Hz, what is the input impedance seen at the Liverpool side? iii) If the characteristic impedance of the transmission line is 20 Ω and the AC signal has a frequency of 10 kHz, what is the input impedance seen at the Liverpool side? iv) Use the results obtained from Questions Q1b-i) to Q1b- to explain why the characteristic impedance is generally a very important parameter at high frequencies, but not so important at low frequencies
Answer #1

we emouo thot impu impedance ot a to time is iven 3AlD -3oo ti20 (6.283ND Am= 20 (4.99くー1.12) J2 20+i 10o (80)Zm= 20 (0.38)(-St. 6チ (iv) ス Zyl 르으 Zn, 쒼를) 辻_ 3